OpenStudy (anonymous):

Find the center and radius of the circle with equation (x+2)²+(y+10)²=4

OpenStudy (yttrium):

An easy one :( C: (-2. -10) r = 2

OpenStudy (unklerhaukus):

A circle with radius \(r\), centred at the origin \[x^2+y^2=r^2\] if instead the circle is centered at (a,b) \[(x-a)^2+(y-b)^2=r^2\]